OpenStudy (anonymous):

Solve of A (a+1)/a + (a+4)/a = 6

OpenStudy (anonymous):

((a+1)+(a+4))/a=6 2a+5=6a 4a=5 a=5/4

OpenStudy (anonymous):

[(a+1+a+4)/a]=6 (2a+5)/a=6 2a+5=6a a=5/4